OpenStudy (anonymous):

Kara's custom tees experienced fixed costs of $300 and variable costs of $2 a shirt. Write an equation that can be used to determine the total expenses encountered by Kara's custom tees. Let x be the number of shirts and C(x) be the cost of producing the shirts The cost of producing 5 shirts is __

OpenStudy (anonymous):

The equation is: C(x)= 200+3x

OpenStudy (anonymous):

just plug in 5 for x

OpenStudy (anonymous):

200+3(5) would I times or add?

OpenStudy (anonymous):

The cost of producing shirts is: C(5)=200+3(5) C(5)=200+15 C(5)=215

OpenStudy (anonymous):

I have a ? for u if you dont mind translate the phrase into a mathematical expression then simplify six times a number minus three times the same number

OpenStudy (anonymous):

6*x-3*x

OpenStudy (anonymous):

then what would it simplify too

OpenStudy (anonymous):

6x-3x you could wright it like this to mybad so then i believe you just subtcract